Show that if n ∣ m, where n and m are integers greater than 1, and if a ≡ b (mod m), where a and b are integers, then a ≡ b (mod n).


1
Expert's answer
2021-08-11T15:34:08-0400

If "a ≡ b (mod\ m)" , where "a" and "b" are integers, then "m|(a-b)," that is "a-b=mt" for some integer "t." Since "n|m,m=nk" for some integer "k." Then "a-b=(nk)t=n(kt)," and hence "n|(a-b)." We conclude that "a ≡ b (mod\ n)."